Solve the following system by substitution. Substitute the expression in the second equation into the first equation and solve.
5x-2y=12
x=-y+1

Answer by stanbon(75887)

5x-2y=12
x=-y+1
------------------
5(-y+1) - 2y = 12
-5y+5 -2y = 12
------
-7y = 7
y = -1
==============
Cheers,
Stan H.
==============